When he is added to the reaction mixture the temperature rises from 39.50 Celsius to 74.2 Celsius calculate the changes in temperature Δ t for this reaction


1
Expert's answer
2021-03-03T02:00:14-0500

T1 = 39.50 ⁰C

T2 = 74.2 ⁰C

"\Delta T" = T2 - T1 = 74.2 - 39.50 = 34.70 ⁰C